Find the vector equation of the line through the point (1, 3) with gradient 2

Grade:12

2 Answers

Arun
4 years ago
Dear Student
 
Vector equation will be
r = (i +3j ) + 2 t (i + 3j)
 
Hope it helps
 
Im case of any query please feel free to ask
 
Regards
Arun
Aditya Gupta
4 years ago
obviously aruns ans is wrong coz when you keep t= – ½ , r becomes zero, but clearly the line y – 3= 2(x – 1) does not pass through the origin.
correct eqn would be:
r= i+3j + t(i+2j)
kindly approve :)